OpenStudy (babystar16):

phil has 50 nickles and dimes worth $4.15. how many dimes does he have?

OpenStudy (anonymous):

\(10x+5(50-x)=415\) if the number of dimes is \(x\)

OpenStudy (anonymous):

Your Answer: Laura. . . it looks like the answer of "daSVgrouch" is incorrect. Here is the correct answer for you: Let n = number of nickels Let d = number of dimes The total of our individual nickels and dimes, is 50. So: n + d = 50 Also, we know that our total in money is 415 cents. 5n + 10d = 415 So, we have 2 equations. n + d = 50 (equation 1) 5n + 10d = 415 (equation 2) We can use addition, subtraction, or substitution. -5n + -5d = -250 (equation 1 multiplied through by "negative 5") 5n + 10d = 415 Now add the two equations. . . 0n + 5d = 165 Simplify. . . 5d = 165 d = 33 If d = 33 and we have 50 total coins, then n = 17. He has 33 dimes. Source(s): I used to teach mathematics in high school.

OpenStudy (anonymous):

dimes are worth ten cents each, and if there are a total of 50 coins and \(x\) dimes then there are \(50-x\) nickles worth 5 cents each. you have a total of 415 cents so set \[10x+5(50-x)=415\]solve for \(x\)

OpenStudy (babystar16):

thank u again im in geomerty and math models in high school right now and i need help badly

OpenStudy (anonymous):

can you solve this equation? or do you still need help with it?